# Efficient End-to-End 6-Dof Grasp Detection Framework for Edge Devices with Hierarchical Heatmaps and Feature Propagation

### v1

@@ Abstract
6-DoF grasp detection is critically important for the advancement of intelligent embodied systems, as it provides feasible robot poses for object grasping. Various methods have been proposed to detect 6-DoF grasps through the extraction of 3D geometric features from RGBD or point cloud data. However, most of these approaches encounter challenges during real robot deployment due to their significant computational demands, which can be particularly problematic for mobile robot platforms, especially those reliant on edge computing devices. This paper presents an Efficient End-to-End Grasp Detection Network (E3GNet) for 6-DoF grasp detection utilizing hierarchical heatmap representations. E3GNet effectively identifies high-quality and diverse grasps in cluttered real-world environments. Benefiting from our end-to-end methodology and efficient network design, our approach surpasses previous methods in model inference efficiency and achieves real-time 6-Dof grasp detection on edge devices. Furthermore, real-world experiments validate the effectiveness of our method, achieving a satisfactory 94% object grasping success rate.
@@ I Introduction
In this work, we propose a novel Efficient End-to-End 6-Dof Grasp detection Network (E3GNet) based on hierarchical heatmap representations, which succeed in detecting high-quality and diverse grasps for real-world object clutters. To the best of our knowledge, our method is the first to achieve real-time 6-Dof grasp detection on edge devices. Our framework first constructs a Global Location Heatmap FPN (feature pyramid network) with a lightweight encoder, Geometry-aware MobileOne, to obtain multi-scale features and predict grasp location heatmaps. Then, the Region Feature Propagation module aggregates graspable region features under the guidance of the location heatmaps. Finally, the graspable region features are fed into a specially designed Rotation Heatmap generation model for grasp rotation detection and refinement, forming scene-level 6-Dof grasps. Experiments on the large-scale grasp dataset show that E3GNet detects more precise and diverse grasps than previous methods on the benchmark. Model inference efficiency experiments on multiple platforms, including edge devices, further prove the computation efficiency of our method. Real-world grasp experiments on the real robot also yield a satisfactory 94% grasp success rate.
@@ III Problem Statement
$\boldsymbol{g}=(x,y,z,\theta,\gamma,\beta,w),$
@@
IV Method
IV-A Global Location Heatmap FPN
@@
$x=\frac{u-c_{x}}{1000\times f_{x}},\quad y=\frac{v-c_{y}}{1000\times f_{y}}$
@@
IV-B Region Feature Propagation
@@
$f_{region}$ $=\text{gridsample}(f_{scene},c+d\times\text{meshgrid}(-\frac{s}{2},\frac{s}{2})),$
@@
IV-C Regional Rotation-Heatmap-based Grasp Detection
@@
V Experiments
V-A Dataset, Metrics and Results
@@
V-B Model Efficiency Experiments
@@ V-C Ablation Studies
V-D Realworld Grasping Result
@@
Table IV presents the performance of our method in real-world environments. Our approach achieves an impressive average grasp success rate of 94 % and successfully clears all 6 scenes, demonstrating its ability to generate high-quality grasps. However, some failures were noted, particularly when the gripper collided with other objects in the clutters. A demo video of real-world grasping can be found at youtube.
VI Conlcusion

### v3

@@ Abstract
6-DoF grasp detection is important for the advancement of intelligent embodied systems, as it provides feasible robot poses for object grasping. Various methods have been proposed to detect 6-DoF grasps through the extraction of 3D geometric features from RGBD or point cloud data. However, most of these approaches encounter challenges during real robot deployment due to their significant computational demands, which can be particularly problematic for mobile robot platforms, especially those reliant on edge computing devices. This paper presents an Efficient End-to-End Grasp Detection Network (E3GNet) for 6-DoF grasp detection utilizing hierarchical heatmap representations. E3GNet effectively identifies high-quality and diverse grasps in cluttered real-world environments. Benefiting from our end-to-end methodology and efficient network design, our approach surpasses previous methods in model inference efficiency and achieves real-time 6-Dof grasp detection on edge devices. Furthermore, real-world experiments validate the effectiveness of our method, achieving a satisfactory 94% object grasping success rate. More details can be found on our project page.
@@ I Introduction
In this work, we propose a novel Efficient End-to-End 6-Dof Grasp detection Network (E3GNet) based on hierarchical heatmap representations, which succeed in detecting high-quality and diverse grasps for real-world object clutters. To the best of our knowledge, our method is the first to achieve real-time 6-Dof grasp detection on edge devices. Our framework first constructs a Global Location Heatmap FPN (feature pyramid network) with a lightweight encoder, Geometry-aware MobileOne [10], to obtain multi-scale features and predict grasp location heatmaps. Then, the Region Feature Propagation module aggregates graspable region features under the guidance of the location heatmaps. Finally, the graspable region features are fed into a specially designed Rotation Heatmap generation model for grasp rotation detection and refinement, forming scene-level 6-Dof grasps. Experiments on the large-scale grasp dataset show that E3GNet detects more precise and diverse grasps than previous methods on the benchmark. Model inference efficiency experiments on multiple platforms, including edge devices, further prove the computation efficiency of our method. Real-world grasp experiments on the real robot also yield a satisfactory 94% grasp success rate.
@@ III-A Problem Statement
$\boldsymbol{g}=(x,y,z,\theta,\gamma,\beta,w),$ (1)
@@
III-B Global Location Heatmap FPN
@@
$x=\frac{u-c_{x}}{1000\times f_{x}},\quad y=\frac{v-c_{y}}{1000\times f_{y}},$ (2)
@@
III-C Region Feature Propagation
@@
$f_{reg}$ $=\text{gridsample}(f_{scene},c+d\times\text{meshgrid}(-\frac{s}{2},\frac{s}{2})),$ (3)
@@
III-D Regional Rotation-Heatmap-based Grasp Detection
@@
IV Experiments
IV-A Dataset, Metrics and Results
@@
IV-B Model Efficiency Experiments
@@ IV-C Ablation Studies
IV-D Realworld Grasping Result
@@
Table IV presents the performance of our method in real-world environments. Our approach achieves an impressive average grasp success rate of 94 % and successfully clears all 6 scenes, demonstrating its ability to generate high-quality grasps. However, some failures were noted, particularly when the gripper collided with other objects in the clutters. Real-world demos and more implementation details can be found on the project page.
V Conlcusion